# CE-GPPO: Coordinating Entropy via Gradient-Preserving Clipping Policy Optimization in Reinforcement Learning

### v2

@@ A.3.2 Derivation
$z^{k+1}_{s,a}-z^{k}_{s,a}$ $=\mathbb{E}_{a^{\prime}\sim\pi_{\theta}(\cdot|s)}\left[\frac{\partial\log\pi_{\theta}(a^{\prime}|s)}{\partial z_{s,a}}\hat{A}(s,a^{\prime})\right]$ (19)
$=\mathbb{E}_{a^{\prime}\sim\pi_{\theta}}\left[(\mathbf{1}_{a^{\prime}=a}-\pi_{\theta}(a|s))\hat{A}(s,a^{\prime})\right]$
$=\sum_{a^{\prime}}\pi_{\theta}(a^{\prime}|s)(\mathbf{1}_{a^{\prime}=a}-\pi_{\theta}(a|s))\hat{A}(s,a^{\prime})$
$=\pi_{\theta}(a|s)\hat{A}(s,a)-\pi_{\theta}(a|s)\sum_{a^{\prime}}\pi_{\theta}(a^{\prime}|s)\hat{A}(s,a^{\prime})$
$=\pi_{\theta}(a|s)\left(\hat{A}(s,a)-\mathbb{E}_{a^{\prime}\sim\pi_{\theta}}[\hat{A}(s,a^{\prime})]\right)$
@@
$\left.\left.\mathbb{E}_{a^{\prime}\sim\pi^{k}}[\log\pi^{k}(a^{\prime}\mid s)]\right)\right]\left[\eta\pi^{k}(a\mid s)\hat{A}(s,a)\right]$
Rewriting this expression as an expectation over the policy:
@@
$\left.\left.\mathbb{E}_{a^{\prime}\sim\pi^{k}}[\log\pi^{k}(a^{\prime}\mid s)]\right)\pi^{k}(a\mid s)\hat{A}(s,a)\right]$
Expanding the expectation:
$\Delta\mathcal{H}\approx$ $-\eta\left(\mathbb{E}_{a\sim\pi^{k}}\left[\log\pi^{k}(a\mid s)\pi^{k}(a\mid s)\hat{A}(s,a)\right]\right$. (22)
$\left.-\mathbb{E}_{a^{\prime}\sim\pi^{k}}[\log\pi^{k}(a^{\prime}\mid s)]\mathbb{E}_{a\sim\pi^{k}}\left[\pi^{k}(a\mid s)\hat{A}(s,a)\right]\right)$
By the baseline-centered advantage assumption, the second expectation term vanishes:
$\mathbb{E}_{a\sim\pi^{k}}\left[\pi^{k}(a\mid s)\hat{A}(s,a)\right]=\mathbb{E}_{a\sim\pi^{k}}\left[\hat{A}(s,a)\right]=0$ (23)
Thus, the expression simplifies to:
@@
$\quad\left.\mathbb{E}_{a^{\prime}\sim\pi^{k}}[\log\pi^{k}(a^{\prime}\mid s)]\cdot 0\right)$
Recognizing the remaining expression as a covariance:
$\Delta\mathcal{H}\approx-\eta\mathrm{Cov}_{a\sim\pi^{k}}\left(\log\pi^{k}(a\mid s),\pi^{k}(a\mid s)\hat{A}(s,a)\right)$ (25)

### v4

@@ A.3.2 Derivation
$z^{k+1}_{s,a}-z^{k}_{s,a}$ $=\eta\mathbb{E}_{a^{\prime}\sim\pi_{\theta}(\cdot|s)}\left[\frac{\partial\log\pi_{\theta}(a^{\prime}|s)}{\partial z_{s,a}}\hat{A}(s,a^{\prime})\right]$ (19)
$=\eta\mathbb{E}_{a^{\prime}\sim\pi_{\theta}}\left[(\mathbf{1}_{a^{\prime}=a}-\pi_{\theta}(a|s))\hat{A}(s,a^{\prime})\right]$
$=\eta\sum_{a^{\prime}}\pi_{\theta}(a^{\prime}|s)(\mathbf{1}_{a^{\prime}=a}-\pi_{\theta}(a|s))\hat{A}(s,a^{\prime})$
$=\eta\pi_{\theta}(a|s)\hat{A}(s,a)-\pi_{\theta}(a|s)\sum_{a^{\prime}}\pi_{\theta}(a^{\prime}|s)\hat{A}(s,a^{\prime})$
$=\eta\pi_{\theta}(a|s)\left(\hat{A}(s,a)-\mathbb{E}_{a^{\prime}\sim\pi_{\theta}}[\hat{A}(s,a^{\prime})]\right)$
@@
$\left.\left.\mathbb{E}_{a^{\prime}\sim\pi^{k}}[\log\pi^{k}(a^{\prime}\mid s)]\right)\right](z^{k+1}_{s,a}-z^{k}_{s,a})$
@@
$\left.\left.\mathbb{E}_{a^{\prime}\sim\pi^{k}}[\log\pi^{k}(a^{\prime}\mid s)]\right)(z^{k+1}_{s,a}-z^{k}_{s,a})\right]$
@@
$\left.\mathbb{E}_{a^{\prime}\sim\pi^{k}}\left[\log\pi^{k}\left(a^{\prime}\mid s\right)\right]\mathbb{E}_{a\sim\pi^{k}}\left[z_{s,a}^{k+1}-z_{s,a}^{k}\right]\right)$
This is precisely the definition of covariance:
$\text{Cov}(X,Y)=\mathbb{E}[XY]-\mathbb{E}[X]\mathbb{E}[Y]$ (21)
where $X=\log\pi^{k}(a\mid s)$, $Y=z_{s,a}^{k+1}-z_{s,a}^{k}$. Therefore:
$\Delta\mathcal{H}\approx-\text{Cov}_{a\sim\pi^{k}}\left(\log\pi^{k}(a\mid s),z_{s,a}^{k+1}-z_{s,a}^{k}\right)$ (22)
Then we substitute the result obtained in step 3 to arrive at the following equation.
$\Delta\mathcal{H}$ $\approx-\operatorname{Cov}_{a\sim\pi^{k}}\left(\log\pi^{k}(a\mid s),\eta\pi^{k}(a\mid s)\hat{A}(s,a)\right)$ (23)
$=-\eta\operatorname{Cov}_{a\sim\pi^{k}}\left(\log\pi^{k}(a\mid s),\pi^{k}(a\mid s)\hat{A}(s,a)\right)$